OpenStudy (anonymous):

Could someone please help me with this? Will fan. A package of modeling clay was shaped like a cone with a radius of 24 cm and a height of 6 cm. Miriam used all the clay to make a cylinder with a radius of 16 cm. What was the height of the cylinder? Use 3.14 to approximate pi, and express your final answer in tenths.

OpenStudy (anonymous):

V of cone = 1/3 * pi * r^2 *h

OpenStudy (anonymous):

now this is equal to the volume of the cylinder

OpenStudy (anonymous):

@--Irwin are u getting it?

OpenStudy (anonymous):

V of cylinder = pi * radius^2 * height equating above two we will get height for this culinder

OpenStudy (anonymous):

erm, I'm sorta getting what your saying

OpenStudy (anonymous):

pi/3*24*24*6 = pi* 16 * 16 * H of cylinder.

OpenStudy (anonymous):

can u solve this? @--Irwin

OpenStudy (anonymous):

yea, one sec lemme work it out

OpenStudy (anonymous):

@- you got it probably

OpenStudy (anonymous):

3617.28 = 803.84 that's without the height because I don't really see how to get it and it isn't in the question. . . .

OpenStudy (anonymous):

see i tell u.

OpenStudy (anonymous):

\[\pi/3*24*24*6 = \pi*16*16*H\] \[1/3*24*24*6 = 16*16*H\] so H = 1/3*24*24*6*1/16*1/16

OpenStudy (anonymous):

H = 4.5 cm

OpenStudy (anonymous):

@--Irwin u got it now.

OpenStudy (anonymous):

ohh, okay yea that makes more sense to me now. Thank you for your help! :)